Homework: Homework #2 Score: 0 of 4 pts Problem 22 Save 5 of 22 (17 complete)> HYV Score: 60%, 18 of 30 pt 1,4 Show Work | Question Help * Consider the following forecast results. Calculate MFE, MAD, and MAPE, using the data for the months January through June. Does the forecast model under-or overforecast? Actual Demand 1,008 972 923 1,059 1.137 991 Month Forecast January February March April May June 1,061 1,042 910 1,014 1,127 1,035 The MFE is(Enter your response rounded to one decimal place and include a minus sign if necessary)
Answer #1
Calculations
n=6
Columns B C D
Actual Forecast Error Abs (Error) Abs (Error)/Actual
1008 1061 -53 53 0.052579365
972 1042 -70 70 0.072016461
923 910 13 13 0.014084507
1059 1014 45 45 0.042492918
1137 1127 10 10 0.008795075
991 1035 -44 44 0.044399596
Sum 6090 6189 -99 235 0.234367922
MAD=Sum(Column C)/n
(235/6) = 39.16667
MAPE=(Sum(Column D)*100)/n
(0.234368*100)/6 = 3.906132
MFE=Sum(Column B)/n
(-99/6) = -16.5
Overforecasted (Since Sum of Forecast is more than sum of actual)